# Stochastic Gradient Descent Jittering for Inverse Problems: Alleviating the Accuracy-Robustness Tradeoff

### v1

@@ Abstract
Inverse problems aim to reconstruct unseen data from corrupted or perturbed measurements. While most work focuses on improving reconstruction quality, generalization accuracy and robustness are equally important, especially for safety-critical applications. Model-based architectures (MBAs), such as loop unrolling methods, are considered more interpretable and achieve better reconstructions. Empirical evidence suggests that MBAs are more robust to perturbations than black-box solvers, but the accuracy-robustness tradeoff in MBAs remains underexplored. In this work, we propose a simple yet effective training scheme for MBAs, called SGD jittering, which injects noise iteration-wise during reconstruction. We theoretically demonstrate that SGD jittering not only generalizes better than the standard mean squared error training but is also more robust to average-case attacks. We validate SGD jittering using denoising toy examples, seismic deconvolution, and single-coil MRI reconstruction. The proposed method achieves cleaner reconstructions for out-of-distribution data and demonstrates enhanced robustness to adversarial attacks.
Introduction
@@
Machine learning (ML) algorithms show significantly better reconstructions than the classical optimization methods. The majority of the work in the field aims to promote reconstruction quality using different deep-learning models, such as transformers (Fabian, Tinaz, and Soltanolkotabi 2022) and diffusion models (Whang et al. 2022). In particular, model-based architectures (MBAs) such as loop unrolling (LU) networks inspired by classical optimization methods, unroll the objective function into a sequence of trainable iterative steps with the forward model $\bm{A}$ involved in every iteration. MBAs are considered to be more interpretable than neural networks that learn a direct inverse without using $\bm{A}$ (Monga, Li, and Eldar 2021), and achieve state-of-the-art reconstructions in different fields, such as natural image restoration (Gilton, Ongie, and Willett 2021) and medical image reconstruction (Bian 2024).
@@
While reconstruction quality is crucial, robustness and generalization accuracy are equally important for IPs. Robustness refers to the model’s ability to handle unwanted noise, while generalization accuracy measures how well the model performs on unseen data. For example, in medical image reconstruction, a model must generalize well to detect subtle details like small bone fractures and be robust enough to eliminate artifacts caused by measurement noise. Despite their importance, robustness and generalization have not been thoroughly explored in the context of IPs, particularly in MBAs. While some studies provide empirical assessments of robustness (Gilton, Ongie, and Willett 2021; Genzel, Macdonald, and März 2022), there is a noticeable lack of theoretical analysis (Gottschling et al. 2020).
@@
In this work, we model MBA training as a bilevel optimization problem without making specific assumptions about neural network architectures. We theoretically analyze its generalization and robustness risks in solving denoising problems and demonstrate that standard mean-squared error (MSE) training may perform poorly in both metrics.
@@
Technologies like adversarial training (AT) (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Kannan, Kurakin, and Goodfellow 2018) and input noise injection (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Kannan, Kurakin, and Goodfellow 2018) are commonly used to enhance robustness, while approaches such as weight perturbation (Wu, Xia, and Wang 2020), stochastic training (Hardt, Recht, and Singer 2016) and sharpness-aware minimization (Foret et al. 2020) are designed to improve generalization. However, there is often a tradeoff between robustness and accuracy (Tsipras et al. 2018; Zhang et al. 2023; Gottschling et al. 2020; Krainovic, Soltanolkotabi, and Heckel 2023). Although strategies like careful loss function design (Pang et al. 2022; Kannan, Kurakin, and Goodfellow 2018; Zhang et al. 2019; Zheng, Zhang, and Mao 2021), data augmentation (Zhang et al. 2017), and robust architectural choices such as dropout (Yang et al. 2020) and teacher-student models (Zhao, Wang, and Wei 2024) have been explored to mitigate this tradeoff in classification tasks, the challenge of balancing accuracy and robustness in IP solvers still remains an open question (Gottschling et al. 2020). Recent work by Krainovic, Soltanolkotabi, and Heckel (2023) introduces jittering in ${\bm{y}}$ for black-box IP neural networks, achieving a comparable level of worst-case robustness to AT but with significant accuracy loss.
@@
  • (network assumptions) We investigate the performance of MBAs with general non-convex neural networks, where training often results in non-unique parameters. This analysis is crucial for demonstrating improvements across various metrics, particularly in robustness and generalization accuracy.
  • (robustness) We theoretically prove that SGD jittering implicitly promotes average-case robustness than regular MSE training. In the experiments, we further demonstrate the effectiveness of SGD jittering under worst-case attacks.
  • (generalization) We are the first to analyze generalization accuracy for IPs with small perturbations in the testing data, rather than relying on PAC-Bayes Theorem, which hypothetically promotes generalization through Hessian regularization.
  • (convergence) We also demonstrate that under the assumption of Lipschitz continuity, SGD jittering converges, ensuring the training is feasible.
Related Work
@@
Smoothness is closely related to the robustness of neural networks. Works like Salman et al. (2019); Cohen, Rosenfeld, and Kolter (2019); Lecuyer et al. (2019); Li et al. (2018) show that regularization in smoothness promotes the robustness of classifiers. On the other hand, the analysis of IP is very different from classifications. Krainovic, Soltanolkotabi, and Heckel (2023) analyzes the worst-case robustness for IPs using a simple linear reconstruction model and directly solves for robustness risk. However, optimizing a convex model results in a solution that excels in only one metric (accuracy or robustness). In this work, we examine nonlinear MBAs in general forms to show improvements in both metrics.
@@
Demonstrating better generalization of classifiers often involves showing larger classification margins (Cao et al. 2019; Kawaguchi, Kaelbling, and Bengio 2017; Lim et al. 2021), but the proof for IPs is very different. A common but indirect approach shows an algorithm can regularize the Hessian of loss landscape with respect to the network parameters. Models that find flat minima in the loss landscape tend to generalize better in practice (Jiang et al. 2019; Keskar et al. 2016; Neyshabur et al. 2017; Foret et al. 2020). Many works demonstrate regularization of the Hessian (Xie, Sato, and Sugiyama 2020; Keskar et al. 2016; Foret et al. 2020) then complete the proof using the PAC-Bayes Theorem in hypothesis. While the PAC-Bayes theorem provides a bound on the generalization error, it does not directly incorporate curvature measures. The exact theoretical mechanism behind it is still not fully understood. In this work, instead of connecting to the PAC-Bayes Theorem, we rewrite the generalization risk and demonstrate that SGD jittering results in a smaller generalization risk than regular MSE training.
@@
Noise injection can be viewed as implicit regularization during training, which is employed to enhance model robustness and generalization. Input injection adds random noises to the signal space, which can be viewed as a data augmentation strategy. It is widely studied in classifications (Fawzi, Moosavi-Dezfooli, and Frossard 2016; Ford et al. 2019; Rusak et al. 2020), and recently shown to promote robustness in IPs (Krainovic, Soltanolkotabi, and Heckel 2023). Weight injection instead perturbed network parameters during training, which is shown to help escape from local minima (Zhu et al. 2018; Nguyen et al. 2019) and promote smooth solutions (Orvieto et al. 2023; Liu et al. 2021; Orvieto et al. 2022). Layer-wise noise injection (Lim et al. 2021) in a recurrent network has been shown to generalize better for classifiers. Weights can also be perturbed before parameter updates to enhance generalization, as in stochastic gradient descent (SGD). We will explore SGD in detail in the next subsection.
@@
SGD training is effective in promoting generalization through implicit regularization (Smith et al. 2021). In large overparameterized models with non-unique global solutions, SGD favors networks with flat minima (Hochreiter and Schmidhuber 1997). The implicit regularization properties of SGD have been explored in contexts such as linear regression (Zou et al. 2021), logistic regression (Ji and Telgarsky 2018), and convolutional networks (Gunasekar et al. 2018). Unlike SGD training, which adds noise to network parameters, the proposed SGD jittering method introduces noise at each iterative update of the reconstructions in the lower-level optimization. Despite its name, SGD jittering is structurally similar to layer-wise noise injection within MBAs.
Definition of Robustness and Generalization
We formally define average-case robustness and generalization risks for a learned inverse mapping $H_{\theta}:{\bm{y}}\mapsto{\bm{x}}$ in IPs. Average-case robustness is also known as the “natural” robustness for average performance (Rice et al. 2021; Hendrycks and Dietterich 2019), which is equally important as in worst-case. During evaluation, a small vector ${\bm{g}}$ sampled from the distribution $\mathcal{P}_{g}$ is added to a data point ${\bm{x}}\in\mathcal{D}$. The new ground-truth becomes ${\bm{x}}_{g}={\bm{x}}+{\bm{g}}$, with the corresponding observation ${\bm{y}}_{g}=\bm{A}{\bm{x}}_{g}+{\bm{z}}={\bm{y}}+\bm{A}{\bm{g}}$.
Definition 1.
@@
$R_{e}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (2)
Definition 2.
@@
$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\in\mathcal{D},{\bm{g}}\in\mathcal{P}_{g}}||{\bm{x}}_{g}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (3)
Robustness measures how close the reconstruction is compared to the data within the dataset, and the generalization accuracy measures how well the reconstruction obeys the physics of the inverse problem even for out-of-distribution (OOD) data. Defining generalization in regression is not straightforward, but for IPs, we can achieve this using the underlying forward models. Typically, a tradeoff between robustness and accuracy exists for classical feedforward neural networks (Zhang et al. 2019; Yang et al. 2020; Stutz, Hein, and Schiele 2019). We aim to illustrate that the SGD jittering technique enhances robustness and accuracy for MBAs without incorporating explicit regularization.
Model-based Architecture Overview
MBAs are inspired by classical optimizations, where the underlying signal ${\bm{x}}$ can be estimated via,
@@
where $r$ is an arbitrary regularization function to ensure the uniqueness of the reconstruction. When $r$ is differentiable, the optimal solution can be solved via gradient descent (GD), iterating through steps $k=1,2,3...$ with a constant step-size $\eta$,
${\bm{x}}_{k+1}={\bm{x}}_{k}-\eta\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})+\eta\nabla r({\bm{x}}_{k}).$
In a loop unrolling (LU) architecture using GD, $\nabla r$ is replaced by a shared-weight neural network, denoted by $f_{\theta}$. The process begins with an initial estimate ${\bm{x}}_{0}$, which is iteratively updated with a finite number of iterations. The final output ${\bm{x}}_{K}$ represents the predicted reconstruction and is compared to the desired target. Different variants of LU are summarized in Monga, Li, and Eldar (2021).
MBA Training Scheme Overview
In this section, we summarize some common training schemes for general regression problems, adapted here for MBA training. Following this, we introduce the proposed SGD jittering method for training MBAs.
@@
MSE training is the most commonly used training scheme for training regression problems and MBAs, where the goal using MBAs is to minimize the following risk w.r.t. $\theta$,
$R(\theta)$ $=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where, }$, (4)
@@
Notice that MBAs do to learn $r$ directly, but instead learn $\nabla r$ using a neural network $f_{\theta}$. We denote $r_{\theta}$ to highlight its dependency of $\theta$. The goal of MSE training is to minimize the squared distance between the output ${\bm{x}}_{K}$ and the ground-truth ${\bm{x}}$. MSE training might overfit the patterns in training data, making it more vulnerable to variations in ${\bm{y}}$.
@@
AT is widely recognized as one of the most effective methods for training robust neural networks (Wang et al. 2021). AT finds the worst-case attack for each training instance and adds it to the input data designed to deceive the model. By learning from these challenging examples, AT improves the model’s robustness. AT aims to minimize the worst-case robustness risk for MBAs as follows,
$R_{\epsilon}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D}}\left[\begin{aligned} &\max_{{\bm{e}}:||{\bm{e}}||_{2}\leq\epsilon}\quad||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\\ &\textrm{s.t. }\hat{{\bm{x}}}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{e}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})\end{aligned}\right].$ (5)
Fast gradient sign method (FGSM) and projected gradient descent (PGD) are methods to find the attack vector ${\bm{e}}$, where (Athalye, Carlini, and Wagner 2018) demonstrate that training a defense model with PGD adversarial attacks is more robust. However, AT is slow as it requires iterative solvers for the attack vector (Shafahi et al. 2019).
AT might suffer from bad generalizations. For example, in noiseless IPs, ${\bm{y}}=\bm{A}{\bm{x}}$, the true inverse $\bm{A}^{-1}$ exists. AT learns an inverse mapping $H_{\theta}$ so that $H_{\theta}({\bm{y}}+{\bm{e}})={\bm{x}}$, but the true estimated should be $\bm{A}^{-1}({\bm{y}}+{\bm{e}})\neq{\bm{x}}$.
@@
Noise injection to the input is widely used in general ML tasks to promote robustness (Salman et al. 2019; Cohen, Rosenfeld, and Kolter 2019; Lecuyer et al. 2019). In IPs that aim to map from ${\bm{y}}$ to ${\bm{x}}$, (Krainovic, Soltanolkotabi, and Heckel 2023) proposes to add zero-mean random Gaussian vector ${\bm{w}}$ in ${\bm{y}}$ to a neural network that learns a direct reconstruction, and demonstrate that with careful selection of the jittering variance, this training scheme can attain a comparable level of worst-case robustness to that achieved by AT. In MBAs, the input jittering approach is expressed as,
$J^{I}_{\sigma_{w}}(\theta)$ $=\mathbb{E}_{{\bm{w}},({\bm{x}},{\bm{y}})\in\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (6)
@@
This method learns an inverse mapping from ${\bm{y}}+{\bm{w}}$ to ${\bm{x}}$, which might suffer from generalization issues in the same way as AT. In fact, Krainovic, Soltanolkotabi, and Heckel (2023) demonstrates a significant accuracy loss using input jittering for black-box neural networks.
Proposed Training Scheme
Another line of work injects random noises layer-wise into the neural network, where (Hodgkinson et al. 2021; Jim, Giles, and Horne 1996; Liu et al. 2020) model the layer-wise injection as stochastic differential equation and show implicit regularization in robustness, and (Lim et al. 2021) added noises to the recurrent neural network and demonstrates better generalization for classification tasks.

### v2

# SGD Jittering: A Training Strategy for Robust and Accurate Model-Based Architectures

@@ Abstract
Inverse problems aim to reconstruct unseen data from corrupted or perturbed measurements. While most work focuses on improving reconstruction quality, generalization accuracy and robustness are equally important, especially for safety-critical applications. Model-based architectures (MBAs), such as loop unrolling methods, are considered more interpretable and achieve better reconstructions. Empirical evidence suggests that MBAs are more robust to perturbations than black-box solvers, but the accuracy-robustness tradeoff in MBAs remains underexplored. In this work, we propose a simple yet effective training scheme for MBAs, called SGD jittering, which injects noise iteration-wise during reconstruction. We theoretically demonstrate that SGD jittering not only generalizes better than the standard mean squared error training but is also more robust to average-case attacks. We validate SGD jittering using denoising toy examples, seismic deconvolution, and single-coil MRI reconstruction. Both SGD jittering and its SPGD extension yield cleaner reconstructions for out-of-distribution data and demonstrates enhanced robustness against adversarial attacks.
Machine Learning, ICML
1 Introduction
@@
Machine learning algorithms show significantly better reconstructions than classical optimization methods. Most of the work in the field aims to promote reconstruction quality using different deep-learning models, such as transformers (Fabian et al., 2022) and diffusion models (Whang et al., 2022). In particular, model-based architectures (MBAs) such as loop unrolling (LU) networks inspired by classical optimization methods, unroll the objective function into a sequence of trainable iterative steps with the forward model $\bm{A}$ involved in every iteration. MBAs are considered to be more interpretable than neural networks that learn a direct inverse without using $\bm{A}$ (Monga et al., 2021), and achieve state-of-the-art reconstructions in different fields, such as natural image restoration (Gilton et al., 2021) and medical image reconstruction (Bian, 2024).
@@
While reconstruction quality is crucial, robustness and generalization accuracy are equally important for IPs. Robustness refers to the model’s ability to handle unwanted noise, while generalization accuracy measures how well the model performs on unseen data. For example, in medical image reconstruction, a model must generalize well to detect subtle details like small bone fractures and be robust enough to eliminate artifacts caused by measurement noise. Despite their importance, robustness and generalization have not been thoroughly explored in the context of IPs, particularly in MBAs. While some studies provide empirical assessments of robustness (Gilton et al., 2021; Genzel et al., 2022), there is a noticeable lack of theoretical analysis (Gottschling et al., 2025).
@@
In this work, we model MBA training as a bilevel optimization problem without making specific assumptions about neural network architectures. We theoretically analyze its generalization and robustness risks in solving denoising problems and demonstrate that standard mean squared error (MSE) training may perform poorly in both metrics.
@@
Technologies like adversarial training (AT) (Fawzi et al., 2016; Gilmer et al., 2019; Kannan et al., 2018) and input noise injection (Fawzi et al., 2016; Gilmer et al., 2019; Kannan et al., 2018) are commonly used to enhance robustness, while approaches such as weight perturbation (Wu et al., 2020), stochastic training (Hardt et al., 2016) and sharpness-aware minimization (Foret et al., 2021) are designed to improve generalization. However, there is often a tradeoff between robustness and accuracy (Tsipras et al., 2019; Zhang et al., 2023; Gottschling et al., 2025; Krainovic et al., 2023). Strategies like careful loss function design (Pang et al., 2022; Kannan et al., 2018; Zhang et al., 2019; Zheng et al., 2021), data augmentation (Zhang et al., 2018), and robust architectural choices such as dropout (Yang et al., 2020) and teacher-student models (Zhao et al., 2024) have been explored to mitigate this tradeoff in classification tasks, but the challenge of balancing accuracy and robustness in IP solvers still remains an open question (Gottschling et al., 2025). Recent work by Krainovic et al. (2023) introduces jittering in ${\bm{y}}$ for black-box IP neural networks, achieving a comparable level of worst-case robustness to AT but with significant accuracy loss.
@@
  • Network Assumptions: We investigate the performance of MBAs with general non-convex neural networks, where training often results in non-unique parameters. This analysis is crucial for demonstrating improvements across various metrics, particularly in robustness and generalization accuracy.
  • Robustness: We theoretically prove that SGD jittering implicitly promotes average-case robustness than regular MSE training. In the experiments, we further demonstrate the effectiveness of SGD jittering under worst-case attacks.
  • Generalization: To the best of our knowledge, this is the first work to formally analyze generalization accuracy in inverse problems under small perturbations in the test data, by reformulating the generalization risk specific to inverse problems. This goes beyond the commonly observed heuristic that a flatter loss landscape implies better generalization.
2 Related Work
Robustness and Generalization for IPs
Several strategies have been proposed to improve robustness in solving inverse problems, including training-time exposure to diverse perturbations such as noise injection and data augmentation (Krainovic et al., 2023; Zhou et al., 2019), and adversarial training (Calivá et al., 2021), enabling models to better handle noise during inference. Other approaches focus on architectural innovations—for example, diffusion models have shown inherent robustness to noise in MRI reconstruction (Güngör et al., 2023), while PINN (Peng et al., 2022) embed domain knowledge to enhance stability without compromising interpretability. To improve generalization in IPs, methods like data augmentation with synthetic perturbations (Guan et al., 2024) and domain adaptation via techniques such as CycleGAN (Zhu et al., 2017) have been used to expand the training distribution and improve adaptability to unseen data. Additionally, incorporating geometric constraints, as in (Jiang et al., 2022) for electrocardiographic image reconstruction, has shown improved generalization by embedding prior knowledge into the learning process. While these approaches typically target either robustness or generalization, achieving both simultaneously remains an open and challenging problem.
@@ 2 Related Work
Smoothness is closely related to the robustness of neural networks. Works like Salman et al. (2019); Cohen et al. (2019); Lecuyer et al. (2019); Li et al. (2018) show that regularization in smoothness promotes the robustness of classifiers. On the other hand, the analysis of IP is very different from classifications. Krainovic et al. (2023) analyzes the worst-case robustness for IPs using a simple linear reconstruction model and directly solves for robustness risk. However, optimizing a convex model results in a solution that excels in only one metric (accuracy or robustness). In this work, we explore nonlinear MBAs in their general forms, demonstrating improvements across both metrics.
@@
Demonstrating better generalization of classifiers often involves showing larger classification margins (Cao et al., 2019; Kawaguchi et al., 2022; Lim et al., 2021), but the proof for IPs is very different. A common but indirect approach shows an algorithm can regularize the Hessian of loss landscape with respect to the network parameters. Models that converge to flat minima in the loss landscape tend to generalize better in practice (Jiang* et al., 2020; Keskar et al., 2017; Neyshabur et al., 2017; Foret et al., 2021), although the precise theoretical mechanism underlying this phenomenon remains unclear. Many studies have shown that regularizing the Hessian of the loss landscape (Xie et al., 2021; Keskar et al., 2017; Foret et al., 2021) leads to flatter minima and have directly linked this to reduced generalization error, but this connection still lacks comprehensive theoretical justification. In this work, rather than focusing on the connection to flat landscapes, we proposed a new generalization risk for IPs, and demonstrate that SGD jittering results in a smaller generalization risk than regular MSE training.
@@
Noise injection serves as an implicit regularization technique during training to enhance model robustness and generalization. When applied to input data, it can be viewed as a data augmentation strategy. This approach has been extensively studied in classification tasks (Fawzi et al., 2016; Gilmer et al., 2019; Rusak et al., 2020) and has recently demonstrated effectiveness in improving robustness for IPs (Krainovic et al., 2023). Layer-wise noise injection (Lim et al., 2021) in a recurrent network has been shown to generalize better for classifiers. Weight injection perturbs network parameters during training, which is shown to help escape from local minima (Zhu et al., 2019; Nguyen et al., 2019) and promote smooth solutions (Orvieto et al., 2023; Liu et al., 2021; Orvieto et al., 2022). Weights can also be perturbed before parameter updates to enhance generalization, as in stochastic gradient descent (SGD). We will explore SGD in detail in the next subsection. In a similar vein, Renaud et al. (2024) adds noise to plug-and-play algorithms during evaluation to enhance in-distribution performance, though it does not address its impact on robustness and generalization accuracy.
@@
SGD training is effective in promoting generalization through implicit regularization (Smith et al., 2021). In large overparameterized models with non-unique global solutions, SGD favors networks with flat minima (Hochreiter & Schmidhuber, 1997). The implicit regularization properties of SGD have been explored in contexts such as linear regression (Zou et al., 2021), logistic regression (Ji & Telgarsky, 2019), and convolutional networks (Gunasekar et al., 2018). Unlike SGD training, which adds noise to network parameters, the proposed SGD jittering method introduces noise at each iterative update of the reconstructions in the lower-level optimization. Despite its name, SGD jittering is structurally similar to layer-wise noise injection within MBAs.
3 Definition of Robustness and Generalization
We formally define average-case robustness and generalization risks for a learned inverse mapping $H_{\theta}:{\bm{y}}\mapsto{\bm{x}}$ in IPs. Average-case robustness is also known as the “natural” robustness for average performance (Rice et al., 2021; Hendrycks & Dietterich, 2019), which is equally important as in worst-case. During evaluation, a small vector ${\bm{g}}$ sampled from the distribution $\mathcal{P}_{g}$ is added to a data point ${\bm{x}}\in\mathcal{D}$. The new ground-truth becomes ${\bm{x}}_{g}={\bm{x}}+{\bm{g}}$, with the corresponding observation ${\bm{y}}_{g}=\bm{A}{\bm{x}}_{g}+{\bm{z}}={\bm{y}}+\bm{A}{\bm{g}}$.
Definition 3.1.
@@ 3 Definition of Robustness and Generalization
$R_{e}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D},{\bm{g}}\sim\mathcal{P}_{g}}||{\bm{x}}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (2)
Definition 3.2.
@@
$\mathcal{G}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D},{\bm{g}}\sim\mathcal{P}_{g}}||{\bm{x}}_{g}-H_{\theta}({\bm{y}}_{g})||^{2}_{2}.$ (3)
Robustness measures how close the reconstruction is compared to the data within the dataset, and the generalization accuracy measures how well the reconstruction obeys the physics of the IP even for out-of-distribution (OOD) data. Defining generalization in regression is not straightforward, but for IPs, we can achieve this using the underlying forward models. Typically, a tradeoff between robustness and accuracy exists for classical feedforward neural networks (Zhang et al., 2019; Yang et al., 2020; Stutz et al., 2019). We aim to illustrate that the SGD jittering technique enhances robustness and accuracy for MBAs without incorporating explicit regularization.
4 Model-based Architecture Overview
MBAs draw inspiration from classical optimization techniques, where the underlying signal ${\bm{x}}$ is estimated via,
@@ 4 Model-based Architecture Overview
where $r$ is an arbitrary regularization function to ensure the uniqueness of the reconstruction. When $r$ is differentiable, the optimal solution can be solved via gradient descent (GD) algorithm, iterating through steps $k=1,2,3...$ with a constant step-size $\eta$,
${\bm{x}}_{k+1}={\bm{x}}_{k}-\eta\bm{A}^{\top}(\bm{A}{\bm{x}}_{k}-{\bm{y}})-\eta\nabla r({\bm{x}}_{k}).$ (4)
In a loop unrolling (LU) architecture using GD, the gradient $\nabla r$ is replaced by a neural network, denoted by $f_{\theta}$. The process begins with an initial estimate ${\bm{x}}_{0}$, which is iteratively updated using (4) with a fixed number of iterations, denoted by $K$. The final output ${\bm{x}}_{K}$ represents the predicted reconstruction and is compared to the desired ground-truth. The weights $\theta$ are then updated through end-to-end backpropagation. Different variants of LU are summarized in Monga et al. (2021).
5 MBA Training Scheme Overview
In this section, we summarize some common training schemes for general regression problems, adapted here for MBA training. Following this, we introduce the proposed SGD jittering method.
@@ 5 MBA Training Scheme Overview
Mean-squared error (MSE) is the most commonly used training loss for regression and IPs. In particular, the goal of training a MBA is to minimize the following risk w.r.t. $\theta$,
$R(\theta)$ $=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where, }$, (5)
@@
Notice that MBAs do not directly learn $r$, but instead learn its gradient $\nabla r$ using a neural network $f_{\theta}$. We denote $r_{\theta}$ to highlight its dependency of $\theta$. The lower-level optimization is implemented using iterative steps, where each iteration corresponding to a single module in a MBA. For LU architecture, with a finite and fixed number of iterations $K$, MSE training minimizes the squared distance between the output ${\bm{x}}_{K}$ and the ground-truth ${\bm{x}}$. However, this training scheme can overfit to patterns in training data, potentially making the model more vulnerable to variations in ${\bm{y}}$.
@@
AT is widely recognized as one of the most effective methods for training robust neural networks (Wang et al., 2019). AT finds the worst-case attack for each training instance and adds it to the input data designed to deceive the model. By learning from these challenging examples, AT improves the model’s robustness. For MBA training, AT aims to minimize the worst-case robustness risk as follows,
$R_{\epsilon}(\theta)=\mathbb{E}_{{\bm{x}},{\bm{y}}\sim\mathcal{D}}\left[\begin{aligned} &\max_{{\bm{e}}:||{\bm{e}}||_{2}\leq\epsilon}\quad||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\\ &\textrm{s.t. }\hat{{\bm{x}}}=\operatorname*{arg\,min}_{x}~{}\frac{1}{2}\|{\bm{y}}+{\bm{e}}-\bm{A}{\bm{x}}\|^{2}_{2}+r_{\theta}({\bm{x}})\end{aligned}\right].$ (6)
Fast gradient sign method and projected gradient descent (ProjGD) are common methods to find the attack vector ${\bm{e}}$, where Athalye et al. (2018) demonstrates that ProjGD adversarial attacks are more robust. However, AT is slow as it requires iterative solvers for the attack vector (Shafahi et al., 2019).
AT might suffer from poor generalizations. Imagine in noiseless IPs where ${\bm{y}}=\bm{A}{\bm{x}}$, and assume the inverse $\bm{A}^{-1}$ exists. The goal of AT is to learn an inverse mapping $H_{\theta}$ such that $H_{\theta}({\bm{y}}+{\bm{e}})$ is close to ${\bm{x}}$, for non-zero vector ${\bm{e}}$. However, the true estimated is $\bm{A}^{-1}({\bm{y}}+{\bm{e}})={\bm{x}}+\bm{A}^{-1}{\bm{e}}$. AT learns to ignore the latter term, thus introducing errors that can degrade generalization accuracy.
@@
Noise injection to the input is widely used in general ML tasks to promote robustness (Salman et al., 2019; Cohen et al., 2019; Lecuyer et al., 2019). In IPs that aim to map from ${\bm{y}}$ to ${\bm{x}}$, Krainovic et al. (2023) proposes adding a zero-mean random Gaussian vector ${\bm{w}}$ to ${\bm{y}}$ as input to a neural network that learns a direct reconstruction. They demonstrate that, with careful selection of the jittering variance, this training scheme can attain a comparable level of worst-case robustness to that achieved by AT. In MBAs, the input jittering approach minimizes,
$J^{I}_{\sigma_{w}}(\theta)$ $=\mathbb{E}_{{\bm{w}},({\bm{x}},{\bm{y}})\sim\mathcal{D}}||{\bm{x}}-\hat{{\bm{x}}}||^{2}_{2}\quad\textrm{ where,}$, (7)
@@
This method learns an inverse mapping from ${\bm{y}}+{\bm{w}}$ to ${\bm{x}}$, which might suffer from generalization issues in the same way as AT. In fact, Krainovic et al. (2023) demonstrates a significant accuracy loss using input jittering for black-box neural networks.
6 Proposed Training Scheme